# Fast, robust and laser-free universal entangling gates for trapped-ion quantum computing

### v2

@@ III.2 Coherent coupling to motional degrees-of-freedom
First, a resonant and not phase-modulated $\pi/2$-pulse prepares the qubit in the eigenstate |$\ket{+}=\frac{1}{\sqrt{2}}\left(\ket{0}+\ket{1}\right)$ of the driving field. Then, the phase-modulated driving field is applied. A final $\pi/2$-pulse with variable phase projects the state vector into the $z-$basis, creating Ramsey fringes.
Since the phase-modulated driving field is continuously applied, the state vector experiences continuous Rabi oscillations, traversing the Bloch sphere. Therefore, to map the coherence, we apply the second Ramsey pulse when the state vector is in a coherent superposition on the Bloch sphere equator. We confirm the required evolution time by a projective measurement along the $z-$axis to ensure the Bloch vector is located in the equatorial plane of the Bloch sphere prior to Ramsey readout. By this, we ensure that we choose interaction times such that the dressed basis coincides with the bare basis, that is, when the rotating frame of Eq. (3) concurs with the lab frame.
Fig. 3 depicts the Ramsey fringe contrast measured as outlined above, for multiple steps in time at $\Omega_{1}=2\pi\times$ 94.8 kHz (blue data points) and $\Omega_{1}=2\pi\times$ 61 kHz (red data points), respectively. In both cases, $\Omega_{2}/\Omega_{1}=\leavevmode\nobreak$0.75. At $\Omega_{1}=2\pi\times$ 94.8 kHz, close to the axial trap frequency, significant qubit-phonon entanglement at about 150 $\mu$s occurs, indicated by the reduction of the qubit’s internal coherence.
The recovery of the Ramsey fringe contrast after $\approx 300\leavevmode\nobreak\ \mu$s shows that the ion’s internal qubit states are disentangled again from the vibrational motion. This is the desired condition after an entangling gate has been completed (details in Sec. IV). For comparison, at $\Omega_{1}=2\pi\times 61\leavevmode\nobreak$kHz where the ion-phonon coupling is weak, a smaller reduction of the fringe contrast occurs (red data points in Fig. 3). For this setting, the qubit-phonon entanglement is reduced as compared to the case when the Rabi frequency $\Omega_{1}$ (which determines the energy splitting between the single dressed states) is set close to the motional frequency (blue data points).
Figure 3: Entanglement and disentanglement of qubit states and motional states of a single ion by application of a phase modulated field (Eq.(2)). For Rabi frequency $\Omega_{1}=2\pi\times 94.8$kHz, close to the secular trap frequency, entanglement is signified by a significant reduction of Ramsey fringe contrast at $t=150\mu$s (blue data points). At $t=300\mu$s, qubit and motion are disentangled again, signified by a recovery of the fringe contrast. For Rabi frequency $\Omega_{1}=2\pi\times 61.0$kHz, qubit-phonon coupling, and, therefore, qubit-phonon entanglement is smaller, resulting in a smaller reduction of fringe contrast (red data points). The dotted lines are added to guide the eye.
@@ IV Two-qubit entangling gate
We now apply two phase-modulated fields, each of which is tuned to the resonance of one of the two ionic qubits.
@@ IV.1 Gate description
The idea to use double-dressed states for two-qubit gates was first proposed in [57]. Our gate extends that scheme using phase modulation of a single driving field as an effective second drive, making it robust to amplitude fluctuations.
@@
Here, $\Omega_{1}^{(j,\text{Phase})}$ is the phase modulation frequency of the $j$-th qubit and $\Omega_{1}^{(j,\text{Amp})}$ is the Rabi frequency determined by the RF field amplitude for qubit $j$. All frequencies are angular frequencies.
We set $\Omega_{1}^{\text{Amp}}=\Omega_{1}^{\text{Phase}}=\Omega_{1}$ for both qubits. Then, following similar transformations as for the one qubit interaction (appendix A), we arrive at the double dressed state Hamiltonian
@@
Here, we define the double-dressed basis operators $F_{z}=S_{x}=-\sigma_{z},\ F_{y}=S_{y}=\sigma_{y},\ F_{x}=-S_{z}=-\sigma_{x}$, and we assume the rotating wave approximation (RWA) like in the single-qubit interaction.
We denote the difference between motional mode frequency and Rabi frequency by $\epsilon=\nu-\Omega_{1}$. The term proportional to $F_{y}^{(j)}$ in Eq. (6) may be interpreted as a driving term in the double-dressed basis, driving sidebands with frequency $\Omega_{1}$. Assuming that
@@
this driving field is far-detuned from the sideband’s transition frequencies at $\Omega_{2}/2\pm\nu$ and the term proportional to $F_{y}$ is negligible. The interaction then simplifies to
@@
which is a Mølmer-Sørensen-type Hamiltonian, that creates the effective interaction
$H_{eff}=-\frac{\left(\eta\nu\right)^{2}}{8\epsilon}F_{z}^{(1)}F_{z}^{(2)},$ (9)
at times $\epsilon t=2\pi n$ for non-zero integer $n$ [47, 48]. Therefore, a maximally entangled state can be achieved by initializing the qubits in the $x-$basis (an equal super-position state) and turning on the phase modulated drive for times
@@
where $k$ is an integer. The shortest gate is achieved for $k=0,n=1$ which results in
@@
In this case, the effective interaction is given by $-\frac{\eta\nu}{8}F_{z}^{(1)}F_{z}^{(2)}$, that is, the effective coupling between the two qubits is directly proportional to the Lamb-Dicke parameter and, therefore, proportional to the static magnetic gradient.
The conditional gate operation described above can be interpreted in terms of its optical counterparts [1, 47, 48, 46]. Let us observe Eq. (3) with $\Omega_{2}=0$. The Hamiltonian is then similar to the Cirac-Zoller interaction Hamiltonian [1], where $\Omega_{1}$ replaces the detuning of the laser drive frequency from the optical qubit resonance and where the carrier transition is eliminated. Therefore, in an analogous way, we can implement an entangling interaction by choosing $\Omega_{1}$ to be close to the motional sidebands. The effective second drive is then used to decouple the motional mode dependent frequency shift turning the interaction into a Mølmer-Sørensen (MS)-type gate.
@@ IV.2 Experimental procedure
In order to realize this gate, the qubits are first optically pumped into state $|00\rangle$. Subsequently, they are both initialized in the super-position state $\ket{++}$ by applying a resonant $\pi/2$-pulse to each of the qubits. Then, the phase-modulated driving fields (Eq. (5)) are applied to both qubits.
For our gate scheme, we require $\Omega_{1}^{\text{Amp}}=\Omega_{1}^{\text{Phase}}\equiv\Omega_{1},$ (optimization over the later condition may result in a longer coherence time [56]). To ensure $\Omega_{1}^{\text{Amp}}=\Omega_{1}^{\text{Phase}}$, RF-amplitudes are calibrated by recording and fitting Rabi oscillations of each individually addressed qubit (appendix C).
The phase modulated field is applied for duration $t$ with the Rabi frequency $\Omega_{1}$ set close to the COM mode frequency, $\nu$: The detuning $\epsilon$ between $\nu$ and $\Omega_{1}$ is set such as to satisfy condition Eq. (11). Here, we use $\Omega_{1}=2\pi\times 94.8$ kHz to fulfill this condition. $\Omega_{2}$, characterizing the phase modulation depth, is set to $\Omega_{2}=2\pi\times 71$ kHz, for double dressing and protecting the qubit’s coherence against RF amplitude fluctuations. For a chosen $\Omega_{1}$, the considerations for choosing $\Omega_{2}$ are described in Sec. IV.1 and guided by numerical simulations, searching for the minimal residual excitation of the ions’ motion.
To reconstruct the density matrix $\rho$, after application of the gate, tomography of the two-qubit state is carried out. For this purpose, the sequence of initial state preparation, subsequent two-qubit entangling gate, and measurement is repeated $M$ times $(200<M<600)$. After preparation of the entangled state, in order to effectively measure in different bare state bases, different sets of local single-qubit rotations are applied before a projective measurement in the $z-$basis takes place [58] (appendix D).
@@ IV.3 Experimental results
A full reconstruction of the two qubit density matrix allows calculating the purity $\mathcal{P}=\text{tr}(\rho^{2})$ as well as the negativity $\mathcal{N}$ of the quantum state as basis-independent measures [59]. Here, the purity of a state is a measure for the degree of residual entanglement between the qubits’ internal atomic state and their external motional state. For qubit and motional states that are perfectly disentangled at the gate time as desired, the purity of the state is unity, while deviations from unity are attributed to residual entanglement with phonons. For a maximally entangled state (such as a Bell state $|\Phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ and $|\Psi^{+}\rangle=(|01\rangle+|10\rangle)/\sqrt{2}$) the negativity is 0.5. The evolution of negativity from zero (the unentangled 2-qubit state) towards the value of 0.5 (that indicates a maximally entangled state) reaching the experimental maximum of $0.48^{+2}_{-6}$, is shown in Fig. 4.
@@
The statistical error of the negativity is calculated as the Gaussian standard deviation of a set of numerically sampled density matrices. Here, we assume that the entries of the reconstructed density matrix are the mean value of a normal distribution having the width of the standard deviation assigned during the reconstruction (appendix E). Due to this reconstruction, the error interval might exceed physical limits. In this case, we truncate the statistical error bars of physical quantities. Similarly, we truncate the error interval of fidelity and purity values based on Gaussian propagation of normal distributions.
@@
An oscillation in the state’s purity from initially 1.0 to $\approx$ 0.6 and back to 1.0 is evident in Fig. 4. This is due to varying qubit-phonon entanglement during the gate evolution time. This oscillation is a direct indication of the phase space trajectory generated by the gate, where at half the gate time, the purity is minimal and at the gate time, the purity is recovered (compare also Fig. 3). This indicates that one closed loop in phase space during the evolution of the gate has been completed [48, 50], realizing the fastest possible two-qubit gate for the parameter set used here (compare section IV.1, Eq. 10, where different gate times are discussed).
@@
Figure 4: Time evolution of the two-qubit entangling gate quantified by the two-qubit state’s negativity (blue data points) and purity (red data points), calculated from a full state reconstruction at interaction times $t$. Here, the negativity increases to $0.48^{+2}_{-6}$ at time $t=313\leavevmode\nobreak\ \mu$s. At this evolution time the Bell state $|\Psi^{+}\rangle$ is created whose experimentally reconstructed density matrix is shown in Fig. 5. The oscillation between a purity of $\approx 0.6$ and 1.0 indicates varying qubit-phonon entanglement during the gate evolution signifying no entanglement between qubits and motional states at the beginning and end of the gate sequence. Here, the achieved purity at the gate time $t=313\leavevmode\nobreak\ \mu$s is $0.99^{+1}_{-3}$. Based on Eq. (5), the gate evolution is simulated using $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, $\Omega_{2}=2\pi\times 71\leavevmode\nobreak$kHz, and a trap frequency of $\nu=2\pi\times 98.08$ kHz. The solid lines show the simulated gate evolution with gate time $t_{\text{g}}^{\text{sim}}=309.5\leavevmode\nobreak\ \mu$s. Discrepancies between experimental data and simulation are caused by fluctuations in experimental parameters (notably the not stabilized secular trap frequency) on a daily basis.
@@
Based on the time evolution measurements of the product-state excitation probability in the bare state basis (appendix F), we select a gate time where the product state excitation probability matches the Bell state $|\Phi^{+}\rangle$, i.e. $P_{|00\rangle}=P_{|11\rangle}=0.5$ and $P_{|01\rangle}=P_{|10\rangle}=0$, or the Bell state $|\Psi^{+}\rangle$, i.e. $P_{|00\rangle}=P_{|11\rangle}=0$ and $P_{|01\rangle}=P_{|10\rangle}=0.5$.
@@
The gate time for generating the Bell state $|\Phi^{+}\rangle$ is $t^{\mathrm{exp}}_{\mathrm{g}}=313\leavevmode\nobreak\ \mathrm{\mathrm{\mu}s}$, while the Bell state $|\Psi^{+}\rangle$ is generated after time ($t^{\mathrm{exp}}_{\mathrm{g}}=310\leavevmode\nobreak\ \mathrm{\mathrm{\mu}s}$). Fig. 5 shows reconstructed density matrices for these Bell states, corrected for state detection errors (appendix G). For $t^{\mathrm{exp}}_{g}=313\leavevmode\nobreak\ \mathrm{\mu}$s, we report measurement outcomes of $\mathcal{N}=0.48^{+2}_{-6}$ and $\mathcal{P}=0.99^{+1}_{-3}$.
@@
Once the density matrix is reconstructed, numerically an optimal set of single qubit rotations is computed to rotate the entangled state into a desired Bell state. While the entangling gate’s parameters are chosen such that the produced state is close to |$\ket{\psi}=\ket{\Phi^{+}}(\ket{\Psi^{+}})$, these numerical single-qubit rotations are applied to infer the resulting state fidelity as
@@
where $R_{j}(\alpha_{j},\beta_{j},\gamma_{j})$ represents the rotation of qubit $j$ around the angles $\alpha_{j},\beta_{j},$ and $\gamma_{j}$. A fidelity of $\mathcal{F}=98^{+2}_{-3}$ w.r.t a maximally entangled $|\Phi^{+}\rangle$ Bell state is achieved. For a gate time of $t^{\mathrm{exp}}_{g}=310\leavevmode\nobreak\ \mathrm{\mu}$s, we report $\mathcal{N}=0.47^{+3}_{-6}$ and $\mathcal{P}=0.98^{+2}_{-3}$, and $\mathcal{F}=97^{+3}_{-3}$ w.r.t a maximally entangled $|\Psi^{+}\rangle$ Bell state. Both Bell states can be selectively generated by choosing appropriate gate times.
Figure 5: Reconstructed density matrices of entangled two-qubit states close to $|\Phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ (a, b) and $|\Psi^{+}\rangle=(|01\rangle+|10\rangle)/\sqrt{2}$ (c,d), using the phase modulated driving field for an interaction time of $310\leavevmode\nobreak\ \mathrm{\mu}$s and $313\leavevmode\nobreak\ \mathrm{\mu}$s, respectively. The resulting states close to |$\ket{\Phi^{+}}$ and |$\ket{\Psi^{+}}$ show a negativity of $0.48^{+2}_{-6}$ and $0.47^{+3}_{-6}$, and a purity value of $0.99^{+1}_{-3}$ and $0.98^{+2}_{-3}$, respectively. The reconstructed density matrices result from 200 measurements in each basis.
The calculated gate time of 309.8 µs, obtained using the aforementioned experimental parameters and the analytical expression given in Eq. 11, is in close agreement with the experimental result. By simulating the system and driving Hamiltonian (Eq. 4 and Eq. 5) after applying only the RWA w.r.t the bare state energies $H_{0}=\sum_{i}\omega_{0}^{i}/2\sigma_{z}^{i}$, we yield the numerically gate time of 309.5 $\mu$s, corresponding to the time when phonon entanglement is minimized. From these simulations we expect a unitary gate fidelity of 99.3 %. The residual infidelity is attributed to the quantum stark shift remaining in the double dressed basis, which is on the order of $(\eta\nu)^{2}/4\Omega_{2}$. Therefore, the infidelity scales as $(\pi\eta\nu/2\Omega_{2})^{2}\sim 0.7\%$ By parameter optimization or pulse shaping methods this can be further reduced. The remaining infidelity in the measured entangling gates is explained by the effect of phonon heating.

### v3

@@ III.2 Coherent coupling to motional degrees of freedom
First, a resonant and not phase-modulated $\pi/2$ pulse prepares the qubit in the eigenstate |$\ket{+}=\frac{1}{\sqrt{2}}\left(\ket{0}+\ket{1}\right)$ of the driving field. Then, the phase-modulated driving field is applied. A final $\pi/2$ pulse with variable phase projects the state vector into the $z$ basis, creating Ramsey fringes.
Since the phase-modulated driving field is continuously applied, the state vector experiences continuous Rabi oscillations, traversing the Bloch sphere. Therefore, to map the coherence, we apply the second Ramsey pulse when the state vector is in a coherent superposition on the Bloch sphere equator. We confirm the required evolution time by a projective measurement along the $z$ axis to ensure the Bloch vector is located in the equatorial plane of the Bloch sphere prior to Ramsey readout. Thus, we ensure that we choose interaction times such that the dressed basis coincides with the bare basis, that is, when the rotating frame of Eq. (3) concurs with the lab frame.
Figure 3 depicts the Ramsey fringe contrast measured as outlined above, for multiple steps in time at $\Omega_{1}=2\pi\times$ 94.8 kHz (blue data points) and $\Omega_{1}=2\pi\times$ 61 kHz (red data points), respectively. In both cases, $\Omega_{2}/\Omega_{1}=\leavevmode\nobreak$0.75. At $\Omega_{1}=2\pi\times$ 94.8 kHz, close to the axial trap frequency, significant qubit-phonon entanglement at about 150 µs occurs, indicated by the reduction of the qubit’s internal coherence.
The recovery of the Ramsey fringe contrast after approximately 300 µs shows that the ion’s internal qubit states are disentangled again from the vibrational motion, which is the desired condition after an entangling gate has been completed (see details in Sec. IV). For comparison, at $\Omega_{1}=2\pi\times 61\leavevmode\nobreak$kHz where the ion-phonon coupling is weak, a smaller reduction of the fringe contrast occurs (red data points in Fig. 3). For this setting, the qubit-phonon entanglement is reduced as compared to the case when the Rabi frequency $\Omega_{1}$ (which determines the energy splitting between the single-dressed states) is set close to the motional frequency (blue data points).
Figure 3: Entanglement and disentanglement of qubit states and motional states of a single ion by application of a phase-modulated field [Eq.(2)]. For Rabi frequency $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, close to the secular trap frequency, entanglement is indicated by a significant reduction of Ramsey fringe contrast at $t=150$ µs (blue data points). At $t=300$ µs, qubit and motion are disentangled again, indicated by a recovery of the fringe contrast. For Rabi frequency $\Omega_{1}=2\pi\times 61.0\leavevmode\nobreak$kHz, qubit-phonon coupling—and, therefore, qubit-phonon entanglement—is smaller, resulting in a smaller reduction of fringe contrast (red data points). The dotted lines are added to guide the eye.
@@ IV Two-qubit entangling gate
We now apply two phase-modulated fields, each of which is tuned to the resonance of one of the two ionic qubits. Before giving a theoretical description of the entangling gate in Sec. IV.1 (and more relevant details in Appendix A), as a guide for the reader, we start this section with a less technical explanation of the gate mechanism.
When driving the bare qubit states of each ion in a two-ion Coulomb crystal with Rabi frequency $\Omega$ on their respective resonance, each of these ions experiences a state-dependent force that oscillates with frequency $\Omega$. These state-dependent forces arise from the fact that the two internal states that make up a qubit experience a differential energy shift in a magnetic gradient field. If the relative phase of the two rf driving fields is stable, then this results in a well-defined coherent displacement of both ions exciting a collective vibrational mode (the c.m. mode in the experiments reported here). Thus, the total wave functions of the ions (internal and motional states) can accumulate a geometric phase that is dependent on the qubit state, and the two qubits can become entangled. Ideally, the motional states should return to their initial state without residual entanglement with the qubit states.
In the absence of an external drive, this residual entanglement is rather small [9]. In this work, however, the Rabi frequency $\Omega$ is close to the trap frequency $\nu$, effectively driving the phonons of the c.m. mode close to resonance and greatly increasing the displacement of the ions, resulting in a much larger geometric phase accumulation in a shorter time, similar to other driven geometric phase gates [50, 48, 49].
The relative phase stability between the two rf fields driving the two qubits at their respective resonance frequencies may not be good enough to create useful entangling gates when the states of the qubits are affected by magnetic noise and / or the rf driving fields suffer from frequency and phase fluctuations at the locations of the ions. The rf-induced errors are strongly suppressed here since (i) both rf drives are generated by the same AWG that is locked to an atomic clock (system oscillator stability: $4.0\times 10^{-13}$) and (ii) the variations in the spatial rf phases at a wavelength around $0.0238\leavevmode\nobreak$m are negligible compared to the indeterminacy of the position of the center of mass of each ion’s spatial wave function which is on the order of nanometers.
Both errors—variations of the qubits’ resonance frequencies and of the rf-frequencies—are removed in the dressed basis [Eq. 3]. However, errors due to amplitude fluctuations of the two rf drives remain, even in the dressed basis. A second driving field is then used to dress the dressed states. Thus, double-dressed states are created that are insensitive to amplitude fluctuations as well. Here, the second driving field is created by simply phase modulating the initial rf drives with frequency $\Omega$; thus it does not introduce amplitude fluctuations on its own.
Importantly, the phase modulation is not only a remedy for amplitude fluctuations; at the same time, it is used to decouple undesired unitary terms in the Hamiltonian [see Eqs. (6)-(8)]. Without the phase-modulated drive, the interaction leading to a conditional quantum gate is of the Jaynes-Cummings type, $\sigma_{+}b+$H.c. This finding results in a trap frequency shift that depends on the qubits’ state, meaning that at the end of the gate, the ions will remain entangled with the phonons. However, with phase modulation, we obtain an interaction between qubits and phonons of the type $\sigma_{z}(b+b^{+})$, the Hamiltonian associated with a MS gate.
@@ IV.1 Gate description
The idea of using double-dressed states for two-qubit gates was first proposed in Ref. [57]. Our gate extends that scheme using phase modulation of a single driving field as an effective second drive, making it robust to amplitude fluctuations.
@@
Here, $\Omega_{1}^{(j,\text{phase})}$ is the phase-modulation frequency of the $j$th qubit and $\Omega_{1}^{(j,\text{amp})}$ is the Rabi frequency determined by the rf field amplitude for qubit $j$. All frequencies are angular frequencies.
We set $\Omega_{1}^{\text{amp}}=\Omega_{1}^{\text{phase}}=\Omega_{1}$ for both qubits. Then, following similar transformations as for the one-qubit interaction (Appendix A), we arrive at the double-dressed-state Hamiltonian
@@
Here, we define the double-dressed-basis operators $F_{z}=S_{x}=-\sigma_{z},\ F_{y}=S_{y}=\sigma_{y},\ F_{x}=-S_{z}=-\sigma_{x}$, and we assume the rotating wave approximation (RWA) like in the single-qubit interaction.
We denote the difference between the motional mode frequency and the Rabi frequency by $\epsilon=\nu-\Omega_{1}$. The term proportional to $F_{y}^{(j)}$ in Eq. (6) may be interpreted as a driving term in the double-dressed basis, driving sidebands with frequency $\Omega_{1}$. Assuming that
@@
this driving field is far detuned from the sideband’s transition frequencies at $\Omega_{2}/2\pm\nu$, and the term proportional to $F_{y}$ is negligible. The interaction then simplifies to
@@
which is a Mølmer-Sørensen-type Hamiltonian that creates the effective interaction
$H_{\text{eff}}=-\frac{\left(\eta\nu\right)^{2}}{8\epsilon}F_{z}^{(1)}F_{z}^{(2)},$ (9)
at times $\epsilon t=2\pi n$ for nonzero integer $n$ [47, 48]. Therefore, a maximally entangled state can be achieved by initializing the qubits in the $x$ basis (an equal superposition state) and turning on the phase-modulated drive for times
@@
where $k$ is an integer. The shortest gate is achieved for $k=0,n=1$, which results in
@@
In this case, the effective interaction is given by $-\frac{\eta\nu}{8}F_{z}^{(1)}F_{z}^{(2)}$; that is, the effective coupling between the two qubits is directly proportional to the Lamb-Dicke parameter and, therefore, proportional to the static magnetic gradient.
The conditional gate operation described above can be interpreted in terms of its optical counterparts [1, 47, 48, 46]. Let us observe Eq. (3) with $\Omega_{2}=0$. The Hamiltonian is then similar to the Jaynes-Cummings interaction Hamiltonian [1], where $\Omega_{1}$ replaces the detuning of the laser drive frequency from the optical qubit resonance and where the carrier transition is eliminated. Therefore, in an analogous way, we can implement an entangling interaction by choosing $\Omega_{1}$ to be close to the motional sidebands. The effective second drive is then used to decouple the motional mode-dependent frequency shift, turning the interaction into a Mølmer-Sørensen (MS)-type gate.
@@ IV.2 Experimental procedure
In order to realize this gate, the qubits are first optically pumped into state $|00\rangle$. Subsequently, they are both initialized in the superposition state $\ket{++}$ by applying a resonant $\pi/2$ pulse to each of the qubits. Then, the phase-modulated driving fields [Eq. (5)] are applied to both qubits.
For our gate scheme, we require $\Omega_{1}^{\text{amp}}=\Omega_{1}^{\text{phase}}\equiv\Omega_{1}$ (optimization over the latter condition may result in a longer coherence time [56]). To ensure $\Omega_{1}^{\text{amp}}=\Omega_{1}^{\text{phase}}$, rf amplitudes are calibrated by recording and fitting Rabi oscillations of each individually addressed qubit (Appendix C).
The phase-modulated field is applied for duration $t$ with the Rabi frequency $\Omega_{1}$ set close to the c.m. mode frequency $\nu$: The detuning $\epsilon$ between $\nu$ and $\Omega_{1}$ is set so as to satisfy condition Eq. (11). Here, we use $\Omega_{1}=2\pi\times 94.8$ kHz to fulfill this condition. Note that $\Omega_{2}$, characterizing the phase-modulation depth, is set to $\Omega_{2}=2\pi\times 71$ kHz, for double dressing and protecting the qubit’s coherence against rf amplitude fluctuations. For a chosen $\Omega_{1}$, the considerations for choosing $\Omega_{2}$ are described in Sec. IV.1 and guided by numerical simulations, searching for the minimal residual excitation of the ions’ motion.
To reconstruct the density matrix $\rho$, after application of the gate, tomography of the two-qubit state is carried out. For this purpose, the sequence of initial state preparation, the subsequent gate of two chevrons, and the measurement are repeated $M$ times $(200<M<600)$. After preparation of the entangled state, in order to effectively measure in different bare state bases, different sets of local single-qubit rotations are applied before a projective measurement in the $z$ basis takes place [58] (Appendix D).
@@ IV.3 Experimental results
A complete reconstruction of the two-qubit density matrix allows us to calculate the purity $\mathcal{P}=\text{tr}(\rho^{2})$ as well as the negativity $\mathcal{N}$ of the quantum state as basis-independent measures [59]. Here, the purity of a state is a measure of the degree of residual entanglement between the qubits’ internal atomic states and their external motional state. For qubit and motional states that are perfectly disentangled at the gate time as desired, the purity of the state is unity, while deviations from unity are attributed to residual entanglement with phonons. For a maximally entangled state [such as a Bell state $|\Phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ and $|\Psi^{+}\rangle=(|01\rangle+|10\rangle)/\sqrt{2}$], the negativity is 0.5. The evolution of negativity from zero (the unentangled two-qubit state) towards the value of 0.5 (which indicates a maximally entangled state) reaching the experimental maximum of $0.48^{+2}_{-6}$ is shown in Fig. 4.
@@
The statistical error of the negativity is calculated as the Gaussian standard deviation of a set of numerically sampled density matrices. Here, we assume that the entries in the reconstructed density matrix are the mean value of a normal distribution that has the width of the standard deviation assigned during reconstruction (Appendix E). Because of this reconstruction, the error interval might exceed physical limits. In this case, we truncate the statistical error bars of physical quantities. Similarly, we truncate the error interval of fidelity and purity values based on Gaussian propagation of normal distributions.
@@
An oscillation in the state’s purity from initially 1.0 to approximately 0.6 and back to 1.0 is evident in Fig. 4. This feature is due to the variable entanglement of qubit phonons during the gate evolution time. This oscillation is a direct indication of the phase-space trajectory generated by the gate, where at half the gate time, the purity is minimal and, at the gate time, the purity is recovered (compare also Fig. 3). This finding indicates that one closed loop in phase space during the evolution of the gate has been completed [48, 50], realizing the fastest possible two-qubit gate for the parameter set used here [compare section IV.1, Eq. (10), where different gate times are discussed].
@@
Figure 4: Time evolution of the two-qubit entangling gate quantified by the two-qubit state’s negativity (blue data points) and purity (red data points), calculated from a full state reconstruction at interaction times $t$. Here, the negativity increases to $0.48^{+2}_{-6}$ at time $t=313$ µs. At this evolution time, the Bell state $|\Psi^{+}\rangle$ is created, whose experimentally reconstructed density matrix is shown in Fig. 5. The oscillation between a purity of approximately $0.6$ and 1.0 indicates varying qubit-phonon entanglement during the gate evolution, signifying no entanglement between qubits and motional states at the beginning and end of the gate sequence. Here, the achieved purity at the gate time $t=313$ µs is $0.99^{+1}_{-3}$. Based on Eq. (5), the gate evolution is simulated using $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, $\Omega_{2}=2\pi\times 71\leavevmode\nobreak$kHz, and a trap frequency of $\nu=2\pi\times 98.08$ kHz. The solid lines show the simulated gate evolution with gate time $t_{\text{g}}^{\text{sim}}=309.5$ µs. Discrepancies between experimental data and simulation are caused by fluctuations in experimental parameters (notably, the unstabilized secular trap frequency) on a daily basis.
@@
Based on the time-evolution measurements of the product state excitation probability in the bare state basis (Appendix F), we select a gate time where the product state excitation probability matches the Bell state $|\Phi^{+}\rangle$, i.e. $P_{|00\rangle}=P_{|11\rangle}=0.5$ and $P_{|01\rangle}=P_{|10\rangle}=0$, or the Bell state $|\Psi^{+}\rangle$, i.e. $P_{|00\rangle}=P_{|11\rangle}=0$ and $P_{|01\rangle}=P_{|10\rangle}=0.5$.
@@
The gate time to generate the Bell state $|\Phi^{+}\rangle$ is $t^{\mathrm{exp}}_{\mathrm{g}}=313$ µs, while the Bell state $|\Psi^{+}\rangle$ is generated after time ($t^{\mathrm{exp}}_{\mathrm{g}}=310$ µs). Figure 5 shows reconstructed density matrices for these Bell states, corrected for state detection errors (Appendix G). For $t^{\mathrm{exp}}_{g}=313$ µs, we report measurement outcomes of $\mathcal{N}=0.48^{+2}_{-6}$ and ${\mathcal{P}=0.99^{+1}_{-3}}$.
@@
Once the density matrix is reconstructed, numerically, an optimal set of single-qubit rotations is computed to rotate the entangled state into a desired Bell state. While the entangling gate’s parameters are chosen such that the produced state is close to |$\ket{\psi}=\ket{\Phi^{+}}(\ket{\Psi^{+}})$, these numerical single-qubit rotations are applied to infer the resulting state fidelity as
@@
where $R_{j}(\alpha_{j},\beta_{j},\gamma_{j})$ represents the rotation of qubit $j$ around the angles $\alpha_{j},\beta_{j},$ and $\gamma_{j}$. A fidelity of ${\mathcal{F}=98^{+2}_{-3}}$with respect to a maximally entangled $|\Phi^{+}\rangle$ Bell state is achieved. For a gate time of $t^{\mathrm{exp}}_{g}=310$ µs, we report ${\mathcal{N}=0.47^{+3}_{-6}}$ and ${\mathcal{P}=0.98^{+2}_{-3}}$, and ${\mathcal{F}=97^{+3}_{-3}}$with respect to a maximally entangled $|\Psi^{+}\rangle$ Bell state. Both Bell states can be generated selectively by choosing appropriate gate times.
Figure 5: Reconstructed density matrices of entangled two-qubit states close to $|\Phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ (a),(b) and $|\Psi^{+}\rangle=(|01\rangle+|10\rangle)/\sqrt{2}$ (c),(d), using the phase-modulated driving field for an interaction time of $310$ µs and $313$ µs, respectively. The resulting states close to |$\ket{\Phi^{+}}$ and |$\ket{\Psi^{+}}$ show negativities of $0.48^{+2}_{-6}$ and $0.47^{+3}_{-6}$ and purity values of $0.99^{+1}_{-3}$ and $0.98^{+2}_{-3}$, respectively. The reconstructed density matrices result from 200 measurements in each basis.
The calculated gate time of 309.8 µs, obtained using the aforementioned experimental parameters and the analytical expression given in Eq. (11), is in close agreement with the experimental result. By simulating the system and driving Hamiltonian [Eqs. (4) and (5)] after applying only the RWA with respect to the bare state energies $H_{0}=\sum_{i}\omega_{0}^{i}/2\sigma_{z}^{i}$, we numerically obtain a gate time of 309.5 µs, corresponding to the time when phonon entanglement is minimized. From these simulations, we expect a unitary gate fidelity of 99.3%. The residual infidelity is attributed to the quantum Stark shift remaining in the double-dressed basis, which is on the order of $(\eta\nu)^{2}/4\Omega_{2}$. Therefore, the infidelity scales as $(\pi\eta\nu/2\Omega_{2})^{2}\sim 0.7\%$ By parameter-optimization or pulse-shaping methods, this value can be further reduced. The remaining infidelity in the measured entangling gates is explained by the effect of phonon heating.